OpenStudy (anonymous):

simplify 2x2(4x2+ 9x – 13)

OpenStudy (bahrom7893):

8x^4+18x^3-26x^2

OpenStudy (anonymous):

2*2(4x^2+9x-13)=4(4x^2+9x-13)=16x^2+36x-52

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Yeah, Thanks (:

OpenStudy (anonymous):

welcome